Challenging Your Students To Do Their Best

Carl C. Cowen Purdue University and the Mathematical Biosciences Institute at Ohio State University

Slides from presentation to Project NExT Workshop at Boulder, Colorado, 28 July 2003

Outline

• Introduction

• Challenging yourself

• Expectations of your students

• Some techniques that work for me

• Discussion

Introduction/Assumptions

• No one knows all the answers (certainly not me!)

• You never stop learning (unless you want to)

• Teaching and learning are individual skills – what works for you is probably not the same as what works for your colleague down the hall

• Nothing works with all your students – you will fail to connect with some of them

• No matter what the level of sophistication of the entering students, we all want them to leave changed, with broader and deeper knowledge and experience

Challenging yourself

• Being a college professor can be hard work

• Be clear about what you want and seek a balance

• To challenge your students, you need to challenge yourself

• Since you can’t do everything, consciously choose what to do

• Find somebody to talk to

Expectations of your students

My most important message: (Most of) Your students will live up to your expectations – so set them high!

Expectations of your students

• What are your expectations of your students?

• High but not out of reach

• Communicate your expectations on day 1 and throughout the term

• Why are your students in your class?

• My students didn’t learn < > in the previous class • Don’t let them off the hook!

• Tell the students what they’ve accomplished – and how they are poised to accomplish more

• Reputation exists and matters

Some techniques

• Seat time

• Complex numbers

• Homework

• Technology and comfort • Reading